Help Me solve this :

1)Walt made an extra $9000 last year from a part-time job. He invested part of the money at a 9% annual rate of interest, and the rest at an 8% annual rate of interest. At the end of the year, he made a total of $770 in interest. How much was invested at 8%?

Answer by sassylilmommie(4)

x + y = 9000
x(.09) + y(.08) = 770
x(.08) + y (.08) = 9000(.08)
.08x + .08y = 720
+ .09x + .08y = 770
.17x = 1490
.17x/17 = 1490/.17
x = 8764.70
8764.70 + y = 9000
8764.70 + y – 8764.70 = 9000 – 8764.70
y = 235.30